# Padding Tone - Revealing the role of the padding tokens in T2I Models

### v1

@@ 3.1 Method
In this case, we conduct the intervention before each attention block to ensure that the attention mechanism fully incorporates the tokens we wish to interpret. We use a full prompt and a “clean pads” prompt, whose representations per diffusion layer are denoted as $E^{(l)}_{\text{full}}$ and $E^{(l)}_{\text{clean}}$, respectively. We explore two directions: first, we replace the prompt tokens with clean pads, using Equation 2. If the images generated from these representations still contain prompt-relevant information, it would suggest that the pads are being utilized. Second, we replace the prompt-contextual pads with the clean pads, as in Equation 3. If the resulting images remain unchanged, this would indicate that the pads are not used for encoding information.
@@
Pads $0.76$$\pm 0.022$ $0.23$$\pm 0.015$
Prompt $0.90$$\pm 0.036$ $0.33$$\pm 0.028$
@@
Full $1.0$$\pm 0.0$ $0.34$$\pm 0.020$
Table 3: Average CLIP scores between images generated (with FLUX) with different IDP interventions and either the full prompt or an image generated from the full prompt. ‘Pad’: prompt-contextual pads; ‘Prompt’: prompt tokens; ‘Clean’: a prompt full of pads, used for comparison; ‘Full’: a prompt with real tokens and pads.
@@ 3.2 Results
Table 3 shows the results of our intervention in the diffusion process. The table shows average CLIP scores of images generated with different IDP interventions, to assess the role of pad and prompt tokens. First, we compute CLIP scores vis-a-vis the full text prompt (Prompt column). As may be expected, images generated from the prompt tokens are similar to the prompt text to the same extent as images generated from the full prompt. Interestingly, images generated from only the pads are much more similar to the text prompt than images generated from clean prompts, indicating that pad tokens are used by the diffusion model to produce images that relate to the prompt.
Next, we compute CLIP scores vis-a-vis images generated from the full prompt (Table 3, Image column). The CLIP score between images generated from full prompts and images generated when using only padding tokens in the diffusion is approximately $76$—significantly higher than the score for randomly generated images from a ‘clean’ padding prompt. This is further evidence that the padding tokens contain visual information closely related to the content of the prompt tokens. However, the CLIP score when using images generated with IDP from the prompt tokens is still higher, suggesting that some information is lost when only using padding tokens in the diffusion model.
Finally, we provide qualitative example in Figure 9 and more examples in Figure 10 (Appendix E), which show that images generated from the prompt-contextual pads with IDP have meaningful semantic information. While images generated solely from prompt tokens typically align with the semantic meaning of the prompt, different visual features are often missing when padding tokens are excluded, while the same features are presented in the padding tokens. It appears that the diffusion model uses padding tokens to create additional visual information, while semantic content remains primarily in the prompt tokens.
Figure 9: Images generated with FLUX from different prompt segments show distinct alignments: prompt tokens produce semantically accurate images, while the visual nuance like ’cozy’ emerges only from the prompt-contextual pad tokens.
@@ 4 Related Work
Special tokens and additional computation
@@ Appendix A Data Creation
  • We randomly selected 50 samples from each of the following categories in the Parti dataset:
    • Fine-grained Detail
    • Imagination
    • Simple Detail
    • Style and Format
    • Complex
    • Linguistic Structures
    • Perspective
    • Quantity
  • For each category, we used the following prompt with GPT-4o: Create an alternative CSV with different prompts of similar style and complexity.
  • For the categories Style and Format, and Simple Detail, we repeated this process twice, generating a total of 100 examples for each.
  • In the end, we obtained 500 prompts overall.
@@ Appendix B Attention Between Image and Text in Different Architectures
To condition the generation process on a textual prompt, T2I models typically employ an attention mechanism. There are two popular methods for achieving this: through cross-attention mechanism, used in models like Stable Diffusion 2 and Stable Diffusion XL, and MM-DiT (Esser et al., 2024) blocks, found in models such as FLUX and Stable Diffusion 3. In the cross-attention mechanism, image patches are projected into query vectors $Q$ while text tokens are projected into key and value vectors $K$ and $V$. Essentially, each image patch draws information from the text tokens based on the attention map $A$:
@@ Appendix C Models
    FLUX utilizes a frozen T5 text encoder and CLIP encoder. A key distinction between FLUX and Stable Diffusion models is that the latter incorporates a transformer architecture with self-attention to both the image and text latent representations in the diffusion process. This allows the diffusion model to modify text representations dynamically during the diffusion.
@@ Appendix D Technical Details
Appendix E Qualitative Examples
@@
Figure 11: Examples of images generated from the same prompts with maximum padding and without padding in Stable Diffusion XL and FLUX. Images generated by Stable Diffusion XL maintain consistent quality, while produced by FLUX without padding often miss key details. For example, given the prompt “a compass beside a feather,” images with padding typically include textured paper with text or a manuscript. In contrast, for the prompt “a boy visiting a zoo,” images generated without padding result in vague animal shapes (first column) or hybrids, such as a mix between a giraffe and a horse (third image). However, adding padding leads to more visually coherent animals.
Appendix F Complementary results
Model Clean Pads Full Pads Prompt
flux-schnell 0.039 0.037 0.036 0.036
ldm 0.033 0.037 0.043 0.042
LLaMA unet 0.034 0.035 0.034 0.041
stable diffusion 2 0.037 0.033 0.037 0.034
stable diffusion 3 0.039 0.035 0.046 0.036
stable diffusion XL 0.023 0.036 0.043 0.039
Table 4: Calculated Standard Deviation of CLIP Scores for each Model and different text encoder interventions.

### v2

@@ 3.1 Method
To investigate the role of padding tokens during the diffusion process, we intervene before each attention block at every diffusion step. More specifically, we use two representations: the fully encoded padded prompt representation from the text encoder and another encoded “clean pads” prompt, whose representations per diffusion layer are denoted as $E^{(l)}_{\text{full}}$ and $E^{(l)}_{\text{clean}}$, respectively. We replace the prompt tokens with clean pads at each diffusion step, before each transformer layer using Equation 2, resulting in an image generated solely through self-attention to the prompt-contextual pads. If the images generated from these representations still contain prompt-relevant information, it would suggest that the pads are being utilized during the diffusion process.
@@
Full $1.0$$\pm 0.0$ $0.33$$\pm 0.037$
Prompt $0.91$$\pm 0.003$ $0.33$$\pm 0.038$
Pads $0.72$$\pm 0.008$ $0.22$$\pm 0.054$
Pads First $0.58$$\pm 0.008$ $0.21$$\pm 0.052$
Pads Second $0.55$$\pm 0.006$ $0.15$$\pm 0.038$
@@
Table 3: Average CLIP scores between images generated (with FLUX) with different IDP interventions and either the full prompt or an image generated from the full prompt. ‘Full’: a prompt with real tokens and pads. ‘Prompt’: prompt tokens; ‘Pad’: prompt-contextual pads; ‘Pad First’: First 20% of the prompt-contextual pads; ‘Pad Second’: Second 20% of the prompt-contextual pads; ‘Clean’: a prompt full of pads, used for comparison;
@@ 3.2 Results
Table 3 shows the results of our intervention in the diffusion process. The table shows average CLIP scores of images generated with different IDP interventions, to assess the role of pad and prompt tokens. First, we compute CLIP scores with regards to the full-text prompt (Prompt column). We find that images generated only from the prompt tokens are similar to the prompt to the same extent as images generated from the full prompt. Images generated only from the prompt-contextual pads are much more similar to the text prompt compared to images generated from clean pads, indicating that pad tokens are used by the diffusion model to represent concepts related to the prompt. That said, the CLIP scores of images generated from the prompt-contextual pads are much lower compared to the CLIP scores of images generated from prompt-contextual pads in LLama-UNet with ITE .
These results are further demonstrated in Figure 10, where we present images generated by  IDP . In the first row, images are generated for the prompt “a cozy living room with a raccoon napping on the couch”. The image generated from the prompt-contextual pads has a cozy style but does not align well with the prompt itself. In the second row, for the prompt “two dogs writing poetry”, the image generated from the prompt-contextual padding tokens retain several visual features present but do not incorporate the concept of writing poetry, which is prominent in the prompt. We hypothesize that in diffusion, pads are used to represent visually related information but do not contain the same semantic information.
To further explore this relationship, we compute CLIP scores with regards to images generated from the full padded prompts (Table 3, Image column). The CLIP score between images generated from full prompts and images generated when using only contextual padding tokens is approximately $72$—significantly higher than the score for randomly generated images from a ‘clean’ padding prompt. This is further evidence that the padding tokens contain visual information closely related to the content of the prompt tokens. Here, similar to the results in Section 2, images generated from the first 20% of the contextual pads contain more information than those generated from the next 20%, indicating that information is not spread evenly and is primarily concentrated in the first contextual pads.
Finally, we provide more qualitative examples in Figure 11 (Appendix G), which show that images generated from the prompt-contextual pads with IDP have meaningful semantic information. While images generated solely from prompt tokens typically align with the semantic meaning of the prompt, different visual features are often missing when padding tokens are excluded, while the same features are presented in the padding tokens. It appears that the diffusion model uses padding tokens to create additional visual information, while semantic content remains primarily in the prompt tokens.
Figure 10: Images generated with FLUX using IDP from different prompt segments show distinct alignments: prompt tokens produce semantically accurate images, while the visual nuance like ’cozy’ emerges only from the prompt-contextual pad tokens.
@@ 4 Related Work
Special tokens and additional computation.
@@ 5 Discussion
Acknowledgments
This research was supported by the Israel Science Foundation (grant 448/20), an Azrieli Foundation Early Career Faculty Fellowship, an AI Alignment grant from Open Philanthropy, and an Academic Gift Award from NVIDIA. HO is supported by the Apple AIML PhD fellowship. The authors would like to thank Mor Ventura for her insightful remarks and valuable suggestions. This research was funded by the European Union (ERC, Control-LM, 101165402). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Data Creation
To construct our dataset, we sampled prompts from the Parti dataset and augmented them using GPT-4o. We randomly selected 50 samples from each of the following eight categories: Fine-grained Detail, Imagination, Simple Detail, Style and Format, Complex, Linguistic Structures, Perspective, and Quantity. For each category, we provided GPT-4o with the prompt: Create an alternative CSV with different prompts of similar style and complexity.
To ensure greater diversity, we repeated this process twice for Style and Format and Simple Detail, generating 100 examples for each. Finally, we manually reviewed all generated prompts to verify their diversity and coherence. In total, our dataset comprises 500 curated prompts.
@@ Appendix B Attention Between Image and Text in Different Architectures
To condition the generation process on a textual prompt, T2I models typically employ an attention mechanism. There are two popular methods for achieving this: through cross-attention mechanism, used in models like Stable Diffusion 2 and Stable Diffusion XL, and Multimodal Diffusion Transformer (MM-DiT(Esser et al., 2024) blocks, found in models such as FLUX and Stable Diffusion 3. In the cross-attention mechanism, image patches are projected into query vectors $Q$ while text tokens are projected into key and value vectors $K$ and $V$. Essentially, each image patch draws information from the text tokens based on the attention map $A$:
@@ Appendix C Models
    FLUX utilizes a frozen T5 text encoder and CLIP encoder. A key distinction between FLUX and Stable Diffusion models is that the latter incorporates a transformer architecture with self-attention to both the image and text latent representations in the diffusion process. This allows the diffusion model to modify text representations dynamically during the diffusion. We use the distilled version of FLUX - FLUX-Schnell.
@@ Appendix D Technical Details
Appendix E Additional Results
Model Full EOS
SDXL $0.34\pm 0.036$ $0.30\pm 0.041$
SD2 $0.33\pm 0.033$ $0.31\pm 0.032$
Table 4: Average CLIP Scores for full representation vs images generated using ITE masking all but the EOS token in Stable Diffusion XL and Strable Diffusion 2 models, both using CLIP as the text encoder.
EOS Token Retains Most Semantic Information.
When the CLIP text encoder is used, we observe a phenomenon where semantic information is concentrated in a special token—the EOS token. During CLIP’s training, this token serves as the primary objective for the text encoder, making its representation align with that of the image CLS token. As a result, semantic information naturally aggregates into this token. In Table 4, we observe that the CLIP score of the EOS token is very high, nearly matching that of the image generated from the full representation. This observation has been noted in previous works, such as Ding et al. (2024).
Appendix F Complementary Results
We have removed the error bars from Figure 5 for clarity. The standard deviations are listed in Table 5.
Model Clean Pads Full Pads Prompt
FLUX $0.039$ $0.037$ $0.036$ $0.036$
LDM $0.033$ $0.037$ $0.043$ $0.042$
LLaMA-U $0.034$ $0.035$ $0.034$ $0.041$
SD 2 $0.037$ $0.033$ $0.037$ $0.034$
SD 3 $0.039$ $0.035$ $0.046$ $0.036$
SDXL $0.023$ $0.036$ $0.043$ $0.039$
Table 5: Calculated standard deviation of CLIP scores for each model and different text encoder interventions. LLaMA-U stands for LLama-UNet. SD stands for Stable Diffusion.
Appendix G Qualitative Examples
@@
Figure 12: Examples of images generated from the same prompts with maximum padding* and without padding in Stable Diffusion XL and FLUX. Images generated by Stable Diffusion XL maintain consistent quality, while produced by FLUX without padding often miss key details. For example, given the prompt “a compass beside a feather,” images with padding typically include textured paper with text or a manuscript. In contrast, for the prompt “a boy visiting a zoo,” images generated without padding result in vague animal shapes (first column) or hybrids, such as a mix between a giraffe and a horse (third image). However, adding padding leads to more visually coherent animals. *Maximum padding length is defined as the default number of padding tokens for each model: 77 for SDXL and 512 for FLUX.